OpenStudy (anonymous):

The sum of two numbers is 66. The larger number is 2 more than 3 times the smaller. What are the two numbers? Earn a medal and fan

OpenStudy (triciaal):

x + y = 66 let y be the larger number y = + 2 + 3x x + 3x + 2 = 66 4x = 64 x = 16 y = 3*16 + 2 = 50 check 16 + 50 = 66